OpenStudy (anonymous):

So I was doing pretty good with these until I came to this one I tried all of my answers but none of them worked so I must have done something wrong Here are the choices A) f(x)=x(x+3)(x+4) B) f(x)=3x(x+3)(x+4) C) f(x)=x(x+5)(x+4) D) f(x)=x(3x+1)(x+4)

OpenStudy (anonymous):

f(x)=x^3 +7x^2 +12x this is the equation I'm having trouble with

OpenStudy (triciaal):

first factor out an x what is the quadratic?

OpenStudy (triciaal):

x^2 + 7x + 12 what are the factors of +12 that add up to +7 we will use +3 and +4 so we get (x +3)(x+4) when we factor the quadratic back to the original x^3 +7x^2 +12x = x(x+3)(x+4)

OpenStudy (anonymous):

Thank you so much @triciaal

OpenStudy (triciaal):

oh you are welcome